OpenStudy (anonymous):

There are two factors of -36 such that one factor is 11 less than half of the other factor. Choose all the pairs of these factors.

OpenStudy (anonymous):

x*y=-36 and x=(1/2)y-11 substitute the second eqn into the first eqn [(1/2)y-11]y=-36 distribute, bring everything to one side and simplify (1/2)y^2-11y+36=0 I'm going to multiply both sides by two to make my factoring easier. y^2-22y+72=0 (y-4)(y-18)=0 y=4, y=18 using one of the original equations above x=(1/2)(4)-11=-9 when y is 4 x=(1/2)(18)-11= -2 when y is 18 so... two pairs of factors satisy the given condition 4 with -9 and 18 with -2